### v1

# Score-Based Metropolis-Hastings Algorithms

Abstract
In this paper, we introduce a new approach for integrating score-based models with the Metropolis-Hastings algorithm. While traditional score-based diffusion models excel in accurately learning the score function from data points, they lack an energy function, making the Metropolis-Hastings adjustment step inaccessible. Consequently, the unadjusted Langevin algorithm is often used for sampling using estimated score functions. The lack of an energy function then prevents the application of the Metropolis-adjusted Langevin algorithm and other Metropolis-Hastings methods, limiting the wealth of other algorithms developed that use acceptance functions. We address this limitation by introducing a new loss function based on the detailed balance condition, allowing the estimation of the Metropolis-Hastings acceptance probabilities given a learned score function. We demonstrate the effectiveness of the proposed method for various scenarios, including sampling from heavy-tail distributions.
1 Introduction
Sampling from probability distributions with only access to samples is a longstanding problem with significant attention devoted to it in the statistics and machine learning communities (Goodfellow et al., 2020; Kingma, 2013; Ho et al., 2020; Dinh et al., 2016; Song et al., 2021). A wealth of literature and research is devoted to studying efficient methods for sampling from densities where only a function proportional to the density is known (Hastings, 1970; Wang et al., 2024). A particular instance of this is the Metropolis-Hastings (MH) method, which provides a condition for accepting or rejecting new samples based on knowledge of the function proportional to the density (Robert et al., 2004). The MH algorithm builds a Markov chain based on a proposal distribution and an acceptance function and its convergence is guaranteed when the process satisfies the detailed balance condition. The success of this framework led to a significant amount of research effort devoted to algorithms that make use of the detailed balance criterion. Various MH algorithms have been introduced, defined by their choice of proposal distributions, such as the Random Walk Metropolis (RW) algorithm (Metropolis et al., 1953), the Metropolis Adjusted Langevin Algorithm (MALA) (Roberts and Tweedie, 1996; Roberts and Stramer, 2002), and the preconditioned Crank-Nicolson (pCN) algorithm (Hairer et al., 2014).
While these methods are theoretically well-grounded and provide strong empirical performance, they require knowledge of the unnormalized density, which may be difficult to obtain when only given samples. In recent years, score-based models have gained prominence due to their ability to learn the score function, i.e., the gradient of the log-density, directly from data samples (Song and Ermon, 2019; Song et al., 2020). Consequently, score-based sampling methods, such as the unadjusted Langevin algorithm, have primarily been employed to leverage the learned score function to sample from the target distribution (Grenander and Miller, 1994; Roberts and Tweedie, 1996).
With only knowledge of the score function, the research devoted to sampling using MH criteria is not used within the context of score-based generative modeling. This work proposes a unifying framework that bridges this gap, enabling the application of MH algorithms using only samples and a learned score function. Our analysis is motivated by two questions:
1. Can the acceptance function of an MH algorithm be estimated solely based on the knowledge of an estimated score function? 2. Does including an MH adjustment step improve sample quality for score-based models?
We answer these questions affirmatively by developing an approach to estimate an acceptance probability function from data and evaluating its effectiveness by comparing to ULA algorithms.
2 Related Work
A significant area of related work begins with estimating a score function for the data and then finding efficient ways of sampling using this estimated score function. While majority of such methods focus on designing sampling schemes that converge to the target distribution faster (e.g. through different numerical integrators), we consider our work as a parallel approach which uses an acceptance function to accept or reject different samples. This work lies at the intersection of the literature on score-based models and the Markov Chain Monte Carlo (MCMC) techniques that use Metropolis-Hastings acceptance functions. We then review recent works on score-based models and Metropolis-Hastings algorithms.
Score-Based Models.
As mentioned within the introduction, many of the empirically successful sampling techniques involve sampling using an estimated score function (Song and Ermon, 2019; Song et al., 2021). These methods are trained using the score loss function by minimizing the Fisher divergence between the true and the predicted score (Hyvärinen, 2005; Vincent, 2011). A significant research effort is concerned with understanding the efficacy of these sampling algorithms, with some papers investigating the asymptotic error of score-based diffusion models, as in (Chen et al., 2023; Zhang et al., 2024). Different sampling schemes have been proposed to improve sampling speed and quality (Dockhorn et al., 2021; Lu et al., 2022; Chen et al., 2024). These methods have relied heavily on various score-based samplers such as ULA or the reverse-time SDE of a diffusion process (Song et al., 2021). Notably, Sjöberg et al. (2023) describe a line integration method to compute the acceptance probabilities.
Metropolis-Hastings Methods.
Metropolis-Hastings (MH) algorithms are a class of MCMC algorithms that use an acceptance function to sample from target distributions (Metropolis et al., 1953; Hastings, 1970). These algorithms generate a sequence of samples by proposing a candidate state from a proposal distribution and accepting or rejecting the candidate based on an acceptance criterion that ensures convergence to the target distribution. The efficiency of MH algorithms heavily depends on the choice of the proposal distribution, and significant research effort has focused on designing efficient proposal mechanisms (Rosenthal et al., 2011; Song et al., 2017; Titsias and Dellaportas, 2019; Davies et al., 2023; Lew et al., 2023).
Recent developments in MH algorithms have explored the use of adaptive and state-dependent proposals to improve convergence rates and reduce autocorrelation in the generated samples (Andrieu and Thoms, 2008; Haario et al., 2001). These adaptive MH methods adjust the proposal distribution during the sampling process to better capture the geometry of the target distribution, leading to faster and more accurate sampling (Roberts and Rosenthal, 2009; Brooks et al., 2011; Hirt et al., 2021; Biron-Lattes et al., 2024).
(a) Original, weights $(0.8,0.2)$.
(b) ULA samples, weights $(0.52,0.48)$.
(c) RW samples, weights $(0.79,0.21)$.
(d) MALA samples, weights $(0.79,0.21)$.
Figure 1: Comparison of sampling methods for a mixture of two Gaussians: (a) Original distribution, (b) ULA with the true scores, (c) RW with the true distribution, and (d) MALA with true scores and distribution. The plots demonstrate the impact of slow mixing between modes, highlighting the challenges in low-density regions. We report the empirical weights of the mixture given by each sampling method.
3 Score-Based Metropolis-Hastings
Let $X\sim p$ be a random variable, with support $\mathrm{supp}(X)=\mathcal{X}\subset\mathbb{R}^{d}$ and $p$ be its probability density function. We assume that $p$ is unknown, and only samples of $X$ are observed. Denote these observed samples by $\{x^{\left(i\right)}\}_{i=1}^{N}$, where $N$ is the number of observed samples.
The MH algorithm involves an acceptance function $a:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$, which defines the probability $a(x^{\prime},x)$ of transitioning from the current $x\in\mathcal{X}$ to a proposal $x^{\prime}\in\mathcal{X}$. The common choice of the acceptance function is given by:
$a(x^{\prime},x)=\min\left\{1,\frac{p(x^{\prime})q(x|x^{\prime})}{p(x)q(x^{\prime}|x)}\right\}$ (1)
where $p(x)$ is the target distribution and $q(x|x^{\prime})$ is the proposal distribution. The MH algorithm proceeds as follows:
  • Initialize: Set $x_{1}\in\mathcal{X}$, number of iterations $T$, and proposal distribution $q(x^{\prime}|x)$.
  • For $t=1,\dots,T$:
    • Propose $x^{\prime}\sim q(x^{\prime}|x_{t})$.
    • Compute $a(x^{\prime},x_{t})$ in (1) as the acceptance probability.
    • Draw $u\sim\text{Uniform}(0,1)$. If $u\leq a(x^{\prime},x_{t})$, accept $x_{t+1}=x^{\prime}$, otherwise set $x_{t+1}=x_{t}$.
Note that for computing $a(x^{\prime},x_{t})$ in (1), a function proportional to $p(x)$ is needed. On the other hand, when sampling from $p(x)$, a common strategy is to estimate the score function $\nabla_{x}\log p(x)$ from data. In this work, we will assume that $\nabla_{x}\log p(x)$ is known through prior knowledge or estimated from the observed samples.
Detailed Balance.
Although the acceptance function is often chosen according to (1), under mild conditions, the sufficient condition for convergence to the target distribution is that the acceptance ratio satisfies the detailed balance condition. The detailed balance condition, also known as the Markov reversibility condition, is defined as:
$\forall x,x^{\prime}\in\mathcal{X},\quad\frac{a(x^{\prime},x)}{a(x,x^{\prime})}=\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}.$ (2)
Acceptance functions satisfying the detailed balance condition allow the Markov chain to have the target distribution as its unique stationary distribution under some mild assumptions. In particular, (1) is a particular solution to this condition. Our goal is then to use the knowledge of the stationary condition in (2) such that our sampling scheme is informed by this regime.
Balance Matching.
To do this, we develop a loss function to estimate a valid acceptance function from only the score function based on the detailed balance condition given in (2). In the following proposition, we present the main observation motivating our method. We denote by $\nabla$ the gradient with respect to $x,x^{\prime}\in\mathbb{R}^{d}$, i.e., $\nabla=\nabla_{x,x^{\prime}}$. Thus, for a function $f$ with values in $\mathbb{R}$, the gradient $\nabla f$ has values in $\mathbb{R}^{2d}$.
Proposition 1.
If the gradients of the log acceptance ratio satisfy the following condition: for every $x,x^{\prime}\in\mathcal{X}$,
$\nabla\log\left(\frac{a(x^{\prime},x)}{a(x,x^{\prime})}\right)=\nabla\log\left(\frac{p(x^{\prime})}{p(x)}\right)+\nabla\log\left(\frac{q(x\mid x^{\prime})}{q(x^{\prime}\mid x)}\right)$ (3)
then the acceptance function $a(x^{\prime},x)$ must satisfy the equality:
$\frac{a(x,x^{\prime})}{a(x^{\prime},x)}=\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}.$
Proof.
Proof is provided in appendix C. ∎
It follows from Proposition 1 that matching the gradients of the left-hand side and right-hand side in (2) is a sufficient condition for verifying the detailed balance condition. In other words, if:
$\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})$
$=\nabla\log p(x^{\prime})-\nabla\log p(x)+\nabla\log q(x\mid x^{\prime})$
$\hskip 115.0pt-\nabla\log q(x^{\prime}\mid x),$
then the detailed balance condition is satisfied.
The acceptance function $a(x^{\prime},x)$ can be modeled using an expressive class of functions, such as neural networks, and this function can be optimized to satisfy the condition in (3). In this setup, the gradients $\nabla\log a(x^{\prime},x)$ and $\nabla\log a(x,x^{\prime})$ correspond to the model’s gradients, which can be efficiently computed via backpropagation when the model is differentiable. The terms $\nabla\log p(x)$ and $\nabla\log p(x^{\prime})$ represent the score functions at $x$ and $x^{\prime}$, which can be approximated with a score model. We can express these score functions as:
$\nabla\log p(x)=\nabla_{x,x^{\prime}}\log p(x)=\left(\nabla_{x}\log p(x),\mathbf{0}\right)^{\top},$
and, similarly,
$\nabla\log p(x^{\prime})=\left(\mathbf{0},\nabla_{x^{\prime}}\log p(x^{\prime})\right)^{\top},$
where $\mathbf{0}\in\mathbb{R}^{d}$ is the zero vector. Finally, we note that the gradients of the proposal distribution $q(x^{\prime}|x)$ are typically known by design when the proposal function is specified.
MH with an Estimated Acceptance Functions.
In cases where the only a score function is available, we propose to estimate the acceptance function with balance matching using the estimated score function so that MH algorithms can be applied. We term this approach Score-Based Metropolis-Hastings. For the scope of this paper, we will instantiate our method on three representative MH algorithms, specified by their proposal functions: the Random Walk Metropolis-Hastings (RW), the Metropolis-Adjusted Langevin Algorithm (MALA), and the Preconditioned Crank-Nicolson (pCN) algorithm. We next briefly review these methods.
(a) Standard Random Walk Sampling.
(b) Score-based RW with true scores.
(c) Score-based RW with learned score.
Figure 2: Comparison of sampling methods from a Gaussian distribution $\mathcal{N}(3,2)$ using three approaches: (a) Standard Random Walk Sampling with the acceptance function defined in (1), (b) Score-based RW with an acceptance network trained using true scores, and (c) Score-based RW with an acceptance network trained using a learned score.
Random Walk Metropolis-Hastings: In the RW algorithm, the proposal function is typically a Gaussian centered at the current state, i.e., $q(x^{\prime}\mid x)=\mathcal{N}\left(x,\sigma^{2}I\right)$, where $\sigma^{2}$ controls the step size of the random walk. This proposal is simple and symmetric, making it easy to implement.
Metropolis Adjusted Langevin Algorithm: MALA uses both the current state and the gradient of the log-probability. The proposal distribution is defined as $q(x^{\prime}\mid x)=\mathcal{N}\left(x+\frac{\varepsilon^{2}}{2}\nabla\log p(x),\varepsilon^{2}I\right)$, where $\varepsilon$ is a step size parameter. This approach leverages the local gradient information to propose moves that are more likely to be accepted.
Preconditioned Crank-Nicolson: The pCN algorithm modifies the standard Crank-Nicolson approach by introducing a preconditioning step. The proposal function for pCN is given by $x^{\prime}=\sqrt{1-\beta^{2}}x+\beta\xi$, where $\xi\sim\mathcal{N}(0,I)$ is a standard Gaussian noise term and $\beta$ controls the step size.
When the acceptance function is learned using balance matching, as defined in (3), we refer to these methods as Score RW, Score MALA, and Score pCN, respectively. While we illustrate our results on these three MH algorithms, we note that the proposed framework is general and can be applied to other proposal functions.
Slow Mixing.
There are two main motivations for our approach. The first is the challenge of applying traditional MH algorithms to generative modeling without requiring knowledge of the unnormalized density, while assuming that the score function can be accurately learned. The second motivation arises from limitations of the unadjusted Langevin algorithm (ULA), which, although commonly used for sampling from score-based models, does not guarantee convergence to the true distribution due to the impact of discretization.
Consequently, by incorporating an MH adjustment step, we address a key limitation observed in the ULA when sampling from distributions with multiple modes, particularly when these modes are separated by regions of low density. Specifically, when the data distribution is a mixture $p(x)=\pi p_{1}(x)+(1-\pi)p_{2}(x)$, where $p_{1}(x)$ and $p_{2}(x)$ are distinct and largely disjoint distributions, the gradient of the log probability density, $\nabla_{x}\log p(x)$, becomes problematic. In regions where $p_{1}(x)$ dominates, the score is driven solely by $p_{1}(x)$, and similarly for $p_{2}(x)$. Consequently, ULA, which relies on these score gradients, fails to correctly sample from the mixture distribution as it does not properly account for the mixing proportions $\pi$ and $1-\pi$. This leads to an incorrect estimation of the relative densities between the modes. The faster mixing time of MALA compared to Langevin has been well-established in the literature as in (Dwivedi et al., 2019; Wu et al., 2022).
To illustrate this, we conduct experiments to generate a mixture of Gaussian distributions with $p_{1}=\mathcal{N}\left((5,5)^{\top},I\right)$ and $p_{2}=\mathcal{N}\left((-5,-5)^{\top},I\right)$, with $\pi=0.8$. As can be observed in Figure 1, ULA produces samples that misrepresent the relative weights of the modes. On the other hand, we observe that the original MH algorithms significantly alleviate this problem, generating samples with mode weights close to the truth. This leads to our Score-Based MH framework, which facilitates the application of MH algorithms when only score functions are available.
(a) $\lambda=0$
(b) $\lambda=0.1$
(c) $\lambda=1.0$
Figure 3: Visualizing the acceptance probabilities for different values of $x^{\prime}$ for a given initial state $x=(0.0,0.0)$. We plot the results for training an acceptance network with three different regularization values $\lambda\in\{0,0.1,1.0\}$
4 Algorithm
4.1 Acceptance Loss Function
Ideal Acceptance Loss. We begin with the ideal scenario where the true score function $\nabla\log p$ is known. In this scenario, we propose the following loss function
$\mathcal{L}_{i}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
$\quad-\left.\nabla\log p(X^{\prime})+\nabla\log p(X)\right.$
$\quad\left.-\nabla\log q(X\mid X^{\prime})+\nabla\log q(X^{\prime}\mid X)\right\|^{2}\bigg{]},$ (4)
with $X\sim p$ and $X^{\prime}\sim q$, and the optimization objective is to learn an acceptance function $a^{*}\in\mathcal{F}$, where $\mathcal{F}$ is a set of hypotheses, such that
$a^{*}=\arg\min_{a\in\mathcal{F}}\mathcal{L}_{i}(a).$
(a) Original samples.
(b) ULA.
(c) Score RW.
(d) Score MALA.
(e) Score pCN.
Figure 4: Comparison of different methods on the Moons dataset.
(a) Original samples.
(b) ULA.
(c) Score RW.
(d) Score MALA
(e) Score pCN.
Figure 5: Comparison of different methods on the Pinwheel dataset.
Practical Acceptance Loss. Typically, only samples from the data-generating process are observed, and the true $\nabla\log p$ is unknown. However, score parameterization through a neural network has proven successful in practice. We assume that we have an approximation of the score function previously estimated through a model $\tilde{s}\left(x\right)\approx\nabla\log p\left(x\right)$. Note that this is often the case in practice where an estimated score function $\tilde{s}$ is used in conjunction with ULA for sampling. In this scenario, the proposed practical loss function is
$\mathcal{L}_{p}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
$\quad-\left.\tilde{s}(X^{\prime})+\tilde{s}(X)-\nabla\log q(X\mid X^{\prime})\right.$
$\hskip 80.0pt\left.+\nabla\log q(X^{\prime}\mid X)\right\|^{2}\bigg{]},$ (5)
with $X\sim p$ and $X^{\prime}\sim q$, and the optimization objective is to learn an acceptance function $\tilde{a}\in\mathcal{F}$ where
$\tilde{a}=\arg\min_{a\in\mathcal{F}}\mathcal{L}_{p}(a).$
Illustrative Example.
We present the sampling results from a Gaussian distribution $\mathcal{N}(3,2)$ using the original Random Walk Metropolis-Hastings Algorithm, as shown in Figure 2(a). Next, we show the results of training with the ideal loss function for the Score-based RW in Figure 2(b), followed by the results using the practical loss function for Score-based RW in Figure 2(c).
4.2 Entropy Regularization
Moreover, our algorithm involves an entropy regularization term alongside the proposed loss function defined in Section 4.1. To illustrate its effectiveness, we present the following result as motivation.
Proposition 2 (Estimated Acceptance Function).
Let $x,x^{\prime}\in\mathcal{X}$, and let $M\geq 1$. Then, the acceptance function defined as:
$a_{M}(x^{\prime},x)=\min\left\{\frac{1}{M},\frac{1}{M}\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}\right\}$
satisfies the detailed balance condition in (2).
Proof.
Proof is provided in the appendix. ∎
It follows from Proposition 2 that there are infinitely many solutions to the detailed balance condition, and the acceptance probabilities can become infinitesimally small. While each of these acceptance functions is a mathematically valid solution to the detailed balance condition, in practice, sampling with these low-acceptance probability functions can lead to always rejecting the proposal, resulting in a lack of convergence within a reasonable timeframe.
Therefore, if we solely optimize the loss functions in (4.1) and (4.1), we risk converging to minima that are legitimate solutions but impractical due to their very low acceptance probabilities. To address this issue, we propose adding an entropy regularization term to our loss function. The modified loss function is given by:
$\mathcal{L}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
$-\tilde{s}(X^{\prime})+\tilde{s}(X)-\nabla\log q(X\mid X^{\prime})$
$\;+\nabla\log q(X^{\prime}\mid X)\|^{2}\bigg{]}+\lambda\,\mathbb{E}\left[H(\log a(X^{\prime},X))\right],$ (6)
where
$H(a(X^{\prime},X))=$ $\log(a(X^{\prime},X))a(X^{\prime},X)$
$+\log(1-a(X^{\prime},X))\left(1-a(X^{\prime},X)\right).$
and $\lambda\geq 0$ is the weight of the regularization term. This regularization term encourages acceptance functions with higher entropy acceptance probabilities, thereby avoiding minima that result in very low acceptance probabilities.
(a) Original samples.
(b) ULA.
(c) Score RW.
(d) Score MALA.
(e) Score pCN.
Figure 6: Comparison of different methods on the S-curve dataset.
(a) Original samples.
(b) ULA.
(c) Score RW.
(d) Score MALA.
(e) Score pCN.
Figure 7: Comparison of different methods on the Swiss Roll dataset.
To illustrate the importance of the regularization term, we present the results using the Moons dataset from scikit-learn (Pedregosa et al., 2011). Based on an estimated score function, we estimate an acceptance function to sample from the Moons dataset using the proposed Score RW algorithm. Training is done for different values of the regularization parameter $\lambda$, and the results are plotted in Figure 3(c). It can observed that the acceptance probability is significantly enhanced with the regularization whereas training without the entropy regularization results in acceptance probability close 00. Although this acceptance probability remains mathematically valid (i.e., it will still converge to the target distribution), the convergence time may become impractically long.
4.3 Implementation
The final loss function used to train the acceptance network in our score MH framework consists of two components: the primary loss term, which enforces the reversibility condition, and an entropy regularization term. The primary loss is designed to minimize the difference between the gradients of the log acceptance probability and the gradients of the log posterior and proposal distributions.
$\mathcal{L}_{\text{emp}}(a)=\frac{1}{N}\sum_{i=1}^{N}l_{i}(a)+\frac{\lambda}{N}\sum_{i=1}^{N}h_{i}(a)$ (7)
with
$l_{i}(a)=\bigg{\|}$ $\nabla\log a\left(x^{\prime(i)},x^{(i)}\right)-\nabla\log a\left(x^{(i)},x^{\prime(i)}\right)$
$-\tilde{s}\left(x^{\prime(i)}\right)+\tilde{s}\left(x^{(i)}\right)$
$-\nabla\log q\left(x^{(i)}\mid x^{\prime(i)}\right)+\nabla\log q\left(x^{\prime(i)}\mid x^{(i)}\right)\bigg{\|}^{2}$
and
$h_{i}(a)$ $=\log a\left(x^{\prime(i)},x^{(i)}\right)\cdot a\left(x^{\prime(i)},x^{(i)}\right)$
$\quad+\log\left(1-a\left(x^{\prime(i)},x^{(i)}\right)\right)\cdot\left(1-a\left(x^{\prime(i)},x^{(i)}\right)\right).$
and where $x^{\prime(i)}=\alpha v^{\prime}+(1-\alpha)\tilde{x}$, with $\tilde{x}\sim p$ and $v^{\prime}\sim q(\cdot\mid\tilde{x})$, where $\alpha\in[0,1]$ controls how much $x^{\prime}$ deviates from the data distribution towards the proposal. In the first epochs of training, we begin with smaller values of $\alpha$ to favor data from the true distribution, and then gradually increase $\alpha$ to estimate the acceptance function around the proposal region. This approach is motivated by the fact that the score function may not be well-estimated outside of the data region, and starting with smaller $\alpha$ helps prevent biasing the training with poorly learned scores.
We present the pseudo-code for the proposed loss function in Algorithm 1 and the training procedure in Algorithm 2. We made several design choices that we observed to be beneficial for learning stable acceptance probabilities:
  • Using residual blocks akin to He et al. (2016).
  • Employing smooth activation functions, such as GeLU or Softplus.
  • Implementing annealed sampling from the proposal distribution, where instead of directly sampling from the target distribution, we start from a weighted combination of the observed data samples and the proposal samples.
Algorithm 1 Acceptance Loss Function
Input: Acceptance net $a(x^{\prime},x)$, Score Net $s$, batch of samples $b_{x}$, batch of proposals $b_{x^{\prime}}$, regularization parameter $\lambda$, gradient clipping threshold $C$.
Step 1: Compute gradients $\nabla\log a(x^{\prime},x)$ and $\nabla\log a(x,x^{\prime})$ using Autograd for every $(x,x^{\prime})\in b_{x}\times b_{x^{\prime}}$.
Step 2: Compute score gradients $s(x)$ and $s(x^{\prime})$ for every $(x,x^{\prime})\in b_{x}\times b_{x^{\prime}}$.
Step 3: Compute proposal gradients $\nabla q(x|x^{\prime})$ and $\nabla q(x^{\prime}|x)$ for every $(x,x^{\prime})\in b_{x}\times b_{x^{\prime}}$.
Step 4: Clip all gradients to avoid instability using threshold $C$.
Step 5: Compute the loss $\mathcal{L}_{\text{emp}}$ as defined in (7).
Output: $\mathcal{L}_{\text{emp}}$
Algorithm 2 Training the Acceptance Network
Input: Acceptance Network $a$, Score Network $s$, dataset $D$, number of epochs $N$, sequence $\{\alpha_{i}\}_{i=1}^{N}\subset[0,1]$ (increasing).
For each epoch $i=1\text{ to }N$:
  1. 1.
    Sample a batch $b_{x}$ and a batch $b_{\tilde{x}}$ from $D$.
  2. 2.
    For each $\tilde{x}\in b_{\tilde{x}}$, sample $v^{\prime}\sim q(\cdot\mid\tilde{x})$ and set $x^{\prime}=\alpha_{i}v^{\prime}+(1-\alpha_{i})\tilde{x}$.
  3. 3.
    Compute the loss as defined in Algorithm 1.
  4. 4.
    Update the Acceptance Network $a$ using the Adam optimizer.
Output: Trained Acceptance Network $a$.
Table 1: Quantitative comparison of methods across the datasets using the following metrics: Wasserstein-1 (W1), Wasserstein-2 (W2), and Maximum Mean Discrepancy (MMD).
Dataset ULA Score RW Score MALA Score pCN
Metric W1 W2 MMD W1 W2 MMD W1 W2 MMD W1 W2 MMD
Moons $0.143$ $0.096$ $0.054$ $0.019$ $0.007$ $0.004$ $\bm{0.018}$ $\bm{0.004}$ $\bm{0.002}$ $0.039$ $0.021$ $0.012$
Pinwheel $0.106$ $0.062$ $0.022$ $\bm{0.012}$ $\bm{0.001}$ $\bm{0.001}$ $0.086$ $0.064$ $0.027$ $0.016$ $0.007$ $0.003$
S-curve $0.082$ $0.055$ $0.016$ $0.050$ $0.015$ $\bm{0.004}$ $\bm{0.046}$ $\bm{0.014}$ $\bm{0.004}$ $0.090$ $0.051$ $0.017$
Swiss Roll $2.881$ $14.409$ $0.811$ $2.270$ $8.470$ $0.400$ $\bm{1.399}$ $\bm{4.575}$ $\bm{0.245}$ $14.167$ $113.824$ $6.903$
(a) W1 vs step size
(b) W2 vs step size
(c) MMD vs step size
Figure 8: Influence of step size on the performance of Score-based MALA vs. ULA, evaluated on the Swiss Roll dataset.
5 Empirical Results
In this section, we present the empirical results on benchmark datasets including (Moons, Pinwheel, S-curve, and Swiss Roll), to demonstrate the validity of our approach. The results are presented in Figures 456, and 7.
Figure 9: Average acceptance ratio as a function of $\lambda$. The average acceptance ratio is calculated by evaluating the acceptance network over a uniform grid of $x^{\prime}$ values, with ranges $[-1.5,2.25]$ for the first dimension and $[-1,1.5]$ for the second dimension, while keeping $x=(0,0)$ fixed. The mean acceptance ratio is then obtained by averaging uniformly over the grid.
Results Analysis. The results in Table 1 illustrate the performance of four sampling methods: ULA, Score-based RW, Score-based MALA, and Score-based pCN, across four datasets using three evaluation metrics: Wasserstein-1 (W1), Wasserstein-2 (W2), and Maximum Mean Discrepancy (MMD).
Overall, the Score RW method consistently achieves lower W1 and W2 values compared to ULA across all datasets, highlighting its efficiency in sample-based inference. This trend is especially noticeable on the Moons and Pinwheel datasets, where the W2 metric significantly outperforms that of the ULA method. For more complex datasets like Swiss Roll, the score MH algorithms show noticeable improvements, with Score RW and Score MALA achieving significantly lower W1 and MMD scores when compared to the other methods.
Stability with respect to Step Size: Score MALA vs. ULA.
The results demonstrate that Score MALA is more robust to variations in step size compared to ULA. As seen in Figure 8, the W1, W2, and MMD metrics consistently show lower sensitivity to changes in step size for Score MALA, indicating its stability across different settings.
Effect of Entropy Regularization.
We study the effect of the entropy regularization term on the Moons dataset. As shown in Figure 9, the relationship between the entropy regularization parameter $\lambda$ and the average acceptance ratio indicates a significant impact on the sampling efficiency. As $\lambda$ increases from 00 to around $4$, the average acceptance ratio rises sharply, suggesting that stronger entropy regularization enhances learning more practical acceptance functions. Beyond $\lambda=4$, the acceptance ratio plateaus. This demonstrates that the performance remains robust for larger values of the entropy regularization term.
6 Discussion
In this work we describe a method for estimating the acceptance function given only knowledge of samples and an estimated score function. We demonstrated the successful estimation of an acceptance function applicable across various MH sampling algorithms. While we empirically evaluate on three particular instances, this work opens the door for further implementations and techniques for sampling using acceptance functions. Additionally, we discussed a particular class of acceptance functions that are amenable for efficient sampling using entropy regularization. The proposed methods provide new avenues for combining MH within the context of score-based generative modeling.
Limitations.
The main limitation of the method is the additional computational cost required to train the acceptance network. Furthermore, more advanced architectures need to be proposed to extend this framework to high-dimensional data. Additionally, further theoretical analysis is needed to provide more appropriate priors on which acceptance function should be chosen.
Acknowledgments
Vahid Tarokh’s work was supported in part by the Air Force Office of Scientific Research under award number FA9550-20-1-0397.
Appendix
In this appendix, we first discuss an alternative approximation of the acceptance function that relies solely on the score functions, derived using the Taylor expansion of the log densities. Next, we consider sampling with score-based MALA for heavy-tailed distributions. Third, we provide the proofs for the theoretical results. Finally, we present additional details of the experiments.
Appendix A Taylor Score-based Metropolis-Hastings
In this section, we propose an approximation of the acceptance function based on the Taylor expansion of the log densities.
For an acceptance function $a(x^{\prime}x)$ as defined in Equation 1, the ratio $r(x^{\prime},x)=\frac{p(x^{\prime})q(x|x^{\prime})}{p(x)q(x^{\prime}|x)}$ can be rewritten as:
$\log\left(r(x^{\prime},x)\right)=\log p(x^{\prime})-\log p(x)+\log q(x\mid x^{\prime})-\log q(x^{\prime}\mid x)$ (8)
Let $\tau=x^{\prime}-x$. Assuming that $\|\tau\|$ is sufficiently small, we perform a second-order Taylor series expansion of the logarithmic terms around the point $x$.
Taylor Expansion of $\log p(x^{\prime})$ Around $x$:
Expanding $\log p(x^{\prime})=\log p(x+\tau)$ using a second-order Taylor series:
$\log p(x+\tau)=\log p(x)+\nabla\log p(x)\cdot\tau+\frac{1}{2}\tau^{\top}\nabla^{2}\log p(x)\,\tau+o(\|\tau\|^{2})$ (9)
Therefore,
$\log p(x^{\prime})-\log p(x)=\nabla\log p(x)\cdot\tau+\frac{1}{2}\tau^{\top}\nabla^{2}\log p(x)\,\tau+o(\|\tau\|^{2})$ (10)
Taylor Expansion of $\log p(x)$ Around $x^{\prime}$:
Similarly, expanding $\log p(x)=\log p(x^{\prime}-\tau)$:
$\log p(x^{\prime}-\tau)=\log p(x^{\prime})-\nabla\log p(x^{\prime})\cdot\tau+\frac{1}{2}\tau^{\top}\nabla^{2}\log p(x^{\prime})\,\tau+o(\|\tau\|^{2})$ (11)
Rearranging gives:
$\log p(x)-\log p(x^{\prime})=-\nabla\log p(x^{\prime})\cdot\tau+\frac{1}{2}\tau^{\top}\nabla^{2}\log p(x^{\prime})\,\tau+o(\|\tau\|^{2})$ (12)
Final Result
Therefore, we have that,
$\log p(x^{\prime})-\log p(x)$ $=\frac{1}{2}\left(\log p(x^{\prime})-\log p(x)\right)+\frac{1}{2}\left(\log p(x^{\prime})-\log p(x)\right)$
$=\frac{1}{2}\left[\nabla\log p(x)\cdot\tau+\nabla\log p(x^{\prime})\cdot\tau\right]$
$\quad+\frac{1}{4}\tau^{\top}\left(\nabla^{2}\log p(x)-\nabla^{2}\log p(x^{\prime})\right)\tau+o(\|\tau\|^{2})$
Simplifying, we obtain:
$\log p(x^{\prime})-\log p(x)=\frac{1}{2}\left(\nabla\log p(x)+\nabla\log p(x^{\prime})\right)\cdot\tau+\frac{1}{4}\tau^{\top}\left(\nabla^{2}\log p(x)-\nabla^{2}\log p(x^{\prime})\right)\tau+o(\|\tau\|^{2})$ (13)
Combining the above results, we have:
$\log r(x^{\prime},x)=\frac{1}{2}\left(\nabla\log p(x)+\nabla\log p(x^{\prime})\right)\cdot\tau+\frac{1}{4}\tau^{\top}\left(\nabla^{2}\log p(x)-\nabla^{2}\log p(x^{\prime})\right)\tau+o(\|\tau\|^{2})+\log q(x\mid x^{\prime})-\log q(x^{\prime}\mid x)$
If the distance between the Hessian is negligible, we can approximate the acceptance function as
$\log r(x^{\prime},x)\approx\frac{1}{2}\left(\nabla\log p(x)+\nabla\log p(x^{\prime})\right)\cdot\tau+\left[\log q(x\mid x^{\prime})-\log q(x^{\prime}\mid x)\right]$
We denote this approximation of the acceptance function as Taylor-1 (Averaging). Figure 10 illustrates how this approximation can provide a better result compared to the first-order and second-order Taylor series from Equation 10, while achieving very similar performance to the original Equation 13, which utilizes the Hessian terms. This is particularly important because computing the Hessian numerically (by backpropagation through a score model) is computationally expensive and unstable.
(a) MMD Comparison
(b) W1 Comparison
(c) W2 Comparison
Figure 10: Comparison of Taylor approximations: Taylor-1 and Taylor-2 methods (with and without averaging) evaluated using MMD, W1, and W2 metrics. The experiments were conducted on the Moons dataset.
Additionally, Figure 11 compare the performance of Taylor-1 (Averaging) to score-based Metropolis-Hastings. We can see that for some range the Taylor approximation provides a very good approximation of the Acceptance function as it yields good sampling quality. However, even though some robustness is observed score-based MALA still outperforms the Taylor approximation.
Figure 11: MMD distance comparison between ULA, Taylor-1 (Averaging), and Score-based MH.
Appendix B Metropolis-Hastings and Generalized Extreme Value Distributions
In this section, we study the effect of sampling using ULA, MALA, and score-based MALA algorithms. We analyze the impact of the adjustment step on distributions with heavy tails and hypothesize that as the tail becomes heavier, the adjustment step becomes increasingly more important. This observation suggests that, to accurately model and generate extreme events using score-based models, incorporating the adjustment step is essential.
Generalized Extreme Value (GEV) Distribution
The Generalized Extreme Value (GEV) distribution is commonly used to model the maxima of datasets and is defined by its probability density function (PDF):
$f(x;\xi,\mu,\sigma)=\begin{cases}\frac{1}{\sigma}\left(1+\xi\frac{x-\mu}{\sigma}\right)^{-1-\frac{1}{\xi}}\exp\left(-\left(1+\xi\frac{x-\mu}{\sigma}\right)^{-\frac{1}{\xi}}\right),&\text{if }\xi\neq 0,\\ \frac{1}{\sigma}\exp\left(-\frac{x-\mu}{\sigma}\right)\exp\left(-\exp\left(-\frac{x-\mu}{\sigma}\right)\right),&\text{if }\xi=0,\end{cases}$
where: - $\mu$ is the location parameter, - $\sigma>0$ is the scale parameter, - $\xi$ is the shape parameter (controls the tail heaviness).
The GEV distribution unifies three types of distributions: Gumbel ($\xi=0$), Fréchet ($\xi>0$), and Weibull ($\xi<0$), making it highly flexible for extreme value analysis.
For the Fréchet distribution, moments of order $k$ exist only if $k<\frac{1}{\xi}$. If $k\geq\frac{1}{\xi}$, the moment diverges, leading to infinite values. This property makes the Fréchet distribution particularly suited for modeling heavy-tailed distributions, as it captures the behavior of distributions with heavy tails and undefined higher-order moments.
Figure 12: Comparison between ULA and MALA sampling methods for GEV$(0,1,\xi)$, highlighting the effect of the adjustment step on heavy-tailed distributions. We fix the step size of both ULA and MALA to $0.1$.
In Figure13, we illustrate that score-based MALA achieves competitive performance compared to the ground truth MALA. While MALA uses the standard acceptance function defined in Equation 1, score-based MALA learns a potentially different acceptance function.
(a) W1 Distance comparison.
(b) MMD comparison.
Figure 13: Comparison between score-based MALA (the score network is trained from data), score-based MALA (True Score) where the acceptance network is trained with the true score function, and the original MALA algorithm. Left: W1 distance. Right: MMD.
Appendix C Proofs
Proof of Proposition 1.
Let $x,x^{\prime}\in\mathcal{X}$. Suppose that the gradients on both sides are equal, i.e.,
$\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})=\nabla\log p(x^{\prime})-\nabla\log p(x)+\nabla\log q(x\mid x^{\prime})-\nabla\log q(x^{\prime}\mid x),$
then integrating this equality gives:
$\log a(x^{\prime},x)-\log a(x,x^{\prime})=\log p(x^{\prime})-\log p(x)+\log q(x\mid x^{\prime})-\log q(x^{\prime}\mid x)+C,$
where $C\in\mathbb{R}$ is a constant. Therefore, the acceptance function $a(x,x^{\prime})$ can be expressed as:
$\frac{a(x^{\prime},x)}{a(x,x^{\prime})}=e^{C}\cdot\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}.$
as $e^{C}$ does not depend on $x$, setting $x=x^{\prime}$, we get that
$1=e^{C}$
Therefore we have that $C=0$, which concludes the proof. ∎
Proof of Proposition 2.
To prove that the acceptance function $a_{M}(x,x^{\prime})$ satisfies the detailed balance condition, we need to show that:
$p(x)q(x^{\prime}\mid x)a_{M}(x^{\prime},x)=p(x^{\prime})q(x\mid x^{\prime})a_{M}(x,x^{\prime}).$
Substituting the definition of the acceptance function $a_{M}$, we consider two cases based on the value of the expression $\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}$. Without loss of generality we consider the case when $\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}\leq 1$, then by definition:
$a_{M}(x^{\prime},x)=\frac{1}{M}\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}.$
Substituting this into the detailed balance equation, we have:
$p(x)q(x^{\prime}\mid x)\cdot\frac{1}{M}\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}=\frac{1}{M}p(x^{\prime})q(x\mid x^{\prime}).$
Similarly, since $\frac{p(x)q(x^{\prime}\mid x)}{p(x^{\prime})q(x\mid x^{\prime})}\leq 1$ in this case, we have:
$a_{M}(x,x^{\prime})=\frac{1}{M}.$
Thus, the right-hand side of the detailed balance condition becomes:
$\frac{1}{M}p(x^{\prime})q(x\mid x^{\prime})$
Therefore, both sides are equal:
$\frac{1}{M}p(x^{\prime})q(x\mid x^{\prime})=\frac{1}{M}p(x^{\prime})q(x\mid x^{\prime}),$
which confirms that the detailed balance condition holds in this case. ∎
Appendix D Experiments Details
In this section, we provide descriptions of the datasets used to generate the empirical results from the main text, and outline the neural network architectures and hyperparameter choices. All presented empirical results were compiled using an NVIDIA RTX 3090 GPU.
D.1 Dataset Descriptions
We use four datasets generated using scikit-learn (Pedregosa et al., 2011) and custom code. The parameters provided ensure reproducibility of the results:
  • Moons: This dataset consists of two interlocking crescent-shaped clusters. We generate $10000$ samples with a noise level of $0.1$ using sklearn.datasets.make_moons(n_samples=10000, noise=0.1).
  • Pinwheel: Data points are arranged in six spiral arms. We generate $10000$ samples with a radial standard deviation of $0.5$, tangential standard deviation of $0.05$, and a rate of $0.25$ (which controls the spread of the arms). The dataset is generated using custom code and the following parameters:
    • num_classes=5,
    • radial_std=0.5,
    • tangential_std=0.05,
    • rate=0.25.
    Algorithm 3 Pinwheel Dataset Generation
    1:Input: Number of samples $n$, number of classes $K$, radial standard deviation $\sigma_{r}$, tangential standard deviation $\sigma_{t}$, rotation rate $\alpha$
    2:Output: Dataset $\mathbf{X}\in\mathbb{R}^{n\times 2}$
    3:Generate random class labels $\mathbf{y}\in\{0,1,\dots,K-1\}$ for $n$ samples
    4:Compute angles $\theta_{k}=\frac{2\pi k}{K}$ for each class $k\in\{0,1,\dots,K-1\}$
    5:Generate radial components $r_{i}\sim\mathcal{N}(1,\sigma_{r}^{2})$ for each sample $i$
    6:Compute tangential noise $\delta_{i}\sim\mathcal{N}(0,\sigma_{t}^{2})$ for each sample $i$
    7:Compute angle for each point: $\phi_{i}=\theta_{y_{i}}+\alpha\cdot r_{i}$
    8:Compute Cartesian coordinates:
    $x_{i}=r_{i}\cdot\cos(\phi_{i})+\delta_{i}$
    $y_{i}=r_{i}\cdot\sin(\phi_{i})+\delta_{i}$
    9:Return the dataset $\mathbf{X}=\{(x_{i},y_{i})\}_{i=1}^{n}$
  • S-curve: This dataset consists of $10000$ points distributed along an ”S”-shaped 3D manifold with added Gaussian noise of $0.1$. It is generated using sklearn.datasets.make_s_curve(n_samples=10000, noise=0.1).
  • Swiss Roll: This dataset contains $10000$ points arranged along a 3D spiral-shaped surface, with Gaussian noise of $0.5$ added. It is generated using sklearn.datasets.make_swiss_roll(n_samples=10000, noise=0.5).
All datasets are converted to PyTorch tensors for further processing. The experiments were performed using an NVIDIA RTX 3090 GPU.
D.2 Neural Network Architectures
We now provide detailed descriptions of several different neural network architectures designed for learning the score function (Score Nets), and the acceptance networks (Acceptance Nets).
Score Nets.
The Score Nets architecture is a simple feed-forward neural network consisting of an input layer, two hidden layers, and an output layer. The input dimension is the same as the output dimension, ensuring that the output matches the shape of the input data. The Softplus activation function is applied after each hidden layer to introduce non-linear transformations. The output layer does not use an activation function, as it directly produces the estimated score of the input.
Acceptance Nets.
The Acceptance Nets architecture is a feed-forward neural network designed to compute the acceptance function for a pair of inputs. The network takes two input vectors, concatenates them, and passes the result through an initial fully connected layer with a customized hidden dimension. Following this, the network consists of three residual blocks, each containing fully connected layers with the same hidden dimension, enhanced by GELU activation functions to improve non-linearity and gradient flow. Finally, the output is passed through a fully connected layer, followed by GELU and a Sigmoid activation to ensure the output is between 0 and 1, representing the acceptance probability.
D.3 Hyperparameters
Score Nets.
We summarize the huyperparameters used to train the Score Nets in Table 2.
Table 2: Score-Net Training Hyperparameters
Dataset Optimizer Learning Rate (LR) Epochs Hidden Dimension
Moons Adam $1\times 10^{-3}$ 5000 64
Pinwheel Adam $5\times 10^{-4}$ 2000 512
S-curve Adam $5\times 10^{-4}$ 2000 512
Swiss Roll Adam $5\times 10^{-4}$ 2000 512
Acceptance Nets.
We provdie below the detailed for training the Acceptance network for the various score-based Metropolis-Hastings algorithms.
Table 3: Acceptance-Net Training Hyperparameters
Dataset Optimizer LR Hidden Dimension Residual Layers Epochs $\lambda$
Moons Adam $5\times 10^{-4}$ 256 3 1000 2
Pinwheel Adam $5\times 10^{-4}$ 256 4 200 2
S-curve Adam $5\times 10^{-4}$ 512 4 200 2
Swiss Roll Adam $5\times 10^{-4}$ 512 4 200 1

### v2

@@
 In this paper, we introduce a new approach for integrating score-based models with the Metropolis-Hastings algorithm.
-While traditional score-based diffusion models excel in accurately learning the score function from data points, they lack an energy function, making the Metropolis-Hastings adjustment step inaccessible.
+Score-based diffusion models have emerged as state-of-the-art methods for generative modeling, achieving remarkable performance in accurately learning the score function from data.
+However, these models lack an explicit energy function, making the Metropolis-Hastings adjustment step inaccessible.
 Consequently, the unadjusted Langevin algorithm is often used for sampling using estimated score functions.
-The lack of an energy function then prevents the application of the Metropolis-adjusted Langevin algorithm and other Metropolis-Hastings methods, limiting the wealth of other algorithms developed that use acceptance functions.
+The lack of an energy function prevents the application of the Metropolis-adjusted Langevin algorithm and other Metropolis-Hastings methods, restricting the use of acceptance-function-based algorithms.
 We address this limitation by introducing a new loss function based on the detailed balance condition, allowing the estimation of the Metropolis-Hastings acceptance probabilities given a learned score function.
-We demonstrate the effectiveness of the proposed method for various scenarios, including sampling from heavy-tail distributions.
+We demonstrate the effectiveness of the proposed method for various scenarios, including sampling from score-based diffusion models and heavy-tail distributions.
 1 Introduction
-Sampling from probability distributions with only access to samples is a longstanding problem with significant attention devoted to it in the statistics and machine learning communities (Goodfellow et al., 2020; Kingma, 2013; Ho et al., 2020; Dinh et al., 2016; Song et al., 2021).
-A wealth of literature and research is devoted to studying efficient methods for sampling from densities where only a function proportional to the density is known (Hastings, 1970; Wang et al., 2024).
-A particular instance of this is the Metropolis-Hastings (MH) method, which provides a condition for accepting or rejecting new samples based on knowledge of the function proportional to the density (Robert et al., 2004).
-The MH algorithm builds a Markov chain based on a proposal distribution and an acceptance function and its convergence is guaranteed when the process satisfies the detailed balance condition.
-The success of this framework led to a significant amount of research effort devoted to algorithms that make use of the detailed balance criterion.
-Various MH algorithms have been introduced, defined by their choice of proposal distributions, such as the Random Walk Metropolis (RW) algorithm (Metropolis et al., 1953), the Metropolis Adjusted Langevin Algorithm (MALA) (Roberts and Tweedie, 1996; Roberts and Stramer, 2002), and the preconditioned Crank-Nicolson (pCN) algorithm (Hairer et al., 2014).
-While these methods are theoretically well-grounded and provide strong empirical performance, they require knowledge of the unnormalized density, which may be difficult to obtain when only given samples.
-In recent years, score-based models have gained prominence due to their ability to learn the score function, i.e., the gradient of the log-density, directly from data samples (Song and Ermon, 2019; Song et al., 2020).
-Consequently, score-based sampling methods, such as the unadjusted Langevin algorithm, have primarily been employed to leverage the learned score function to sample from the target distribution (Grenander and Miller, 1994; Roberts and Tweedie, 1996).
-With only knowledge of the score function, the research devoted to sampling using MH criteria is not used within the context of score-based generative modeling.
+Sampling from probability distributions with access only to samples is a longstanding challenge that has received significant attention in machine learning and statistics (Goodfellow et al., 2020; Kingma, 2013; Ho et al., 2020; Dinh et al., 2016; Song et al., 2021).
+In particular, score-based models have emerged as a powerful class of generative algorithms, underpinning state-of-the-art (SOTA) models such as DALL·E 2 and Stable Diffusion (Rombach et al., 2022).
+Given samples from a target distribution, these models first estimate the score function —the gradient of the log-density of the data distribution (Hyvärinen, 2005)—using techniques like sliced score matching (Song et al., 2020) and denoising score matching (Vincent, 2011; Song and Ermon, 2019; Reu et al., ).
+Subsequently, they leverage the learned score function and score-based sampling methods, primarily the Unadjusted Langevin Algorithm (ULA), to sample from the target distribution (Grenander and Miller, 1994; Roberts and Tweedie, 1996).
+In parallel, a wealth of literature focuses on sampling from densities known only up to a normalizing constant (Hastings, 1970; Wang et al., 2024).
+A particular instance of this is the Metropolis-Hastings (MH) method, which provides a condition for accepting or rejecting new samples based on knowledge of a function proportional to the density (Robert et al., 2004).
+The MH method construct a Markov chain based on a proposal distribution and an acceptance function, ensuring convergence when the detailed balance condition is satisfied.
+Various MH algorithms have been introduced, defined by their choice of proposal distributions, such as the Random Walk Metropolis (RW) algorithm (Metropolis et al., 1953), the Metropolis-adjusted Langevin algorithm (MALA) (Roberts and Tweedie, 1996; Roberts and Stramer, 2002), and the preconditioned Crank-Nicolson (pCN) algorithm (Hairer et al., 2014).
+While these methods are theoretically well-grounded, they require knowledge of the unnormalized density, making them impractical when only samples or an estimated score function are available.
+A natural observation arises when considering score-based modeling: despite the extensive research on MH sampling, its integration with score-based methods remains largely unexplored.
+The key technical challenge lies in the absence of an unnormalized density when only the score function (or its estimation) is available.
 This work proposes a unifying framework that bridges this gap, enabling the application of MH algorithms using only samples and a learned score function.
-Our analysis is motivated by two questions:
-1.
-Can the acceptance function of an MH algorithm be estimated solely based on the knowledge of an estimated score function?
-2.
+Consequently, this framework can leverage the strengths of both MH and score-based methods.
+Our analysis is motivated by two key questions:
+A).
+Can we estimate the acceptance function of an MH algorithm solely from an (estimated) score function and samples?
+B).
 Does including an MH adjustment step improve sample quality for score-based models?
-We answer these questions affirmatively by developing an approach to estimate an acceptance probability function from data and evaluating its effectiveness by comparing to ULA algorithms.
+We answer these questions affirmatively by developing an approach to estimate an acceptance function from data. Our contributions are as follows:
+  • We introduce Score Balance Matching, a method for estimating the acceptance function in the MH algorithm using only an estimated score function.
+  • We use the learned acceptance function to incorporate MH steps into score-based samplers like ULA, enhancing the quality and diversity of generated samples.
+  • We evaluate our method’s effectiveness and its robustness to hyperparameters by comparing it to ULA across diverse scenarios.
 2 Related Work
-A significant area of related work begins with estimating a score function for the data and then finding efficient ways of sampling using this estimated score function.
-While majority of such methods focus on designing sampling schemes that converge to the target distribution faster (e.g.
-through different numerical integrators), we consider our work as a parallel approach which uses an acceptance function to accept or reject different samples.
-This work lies at the intersection of the literature on score-based models and the Markov Chain Monte Carlo (MCMC) techniques that use Metropolis-Hastings acceptance functions.
-We then review recent works on score-based models and Metropolis-Hastings algorithms.
+A significant area of related work focuses on estimating a score function and developing efficient sampling strategies based on it.
+While most approaches aim to accelerate convergence to the target distribution (e.g.
+through improved numerical integrators), we consider our work as a parallel approach, incorporating an acceptance function to refine samples.
+This work lies at the intersection of the literature on score-based models and Metropolis-Hastings methods.
+We first review recent advances in both areas.
 Score-Based Models.
-As mentioned within the introduction, many of the empirically successful sampling techniques involve sampling using an estimated score function (Song and Ermon, 2019; Song et al., 2021).
-These methods are trained using the score loss function by minimizing the Fisher divergence between the true and the predicted score (Hyvärinen, 2005; Vincent, 2011).
+Many successful generative models sample from an estimated score function (Song and Ermon, 2019; Song et al., 2021).
+These methods are trained using by minimizing the Fisher divergence between the true and predicted score (Hyvärinen, 2005).
 A significant research effort is concerned with understanding the efficacy of these sampling algorithms, with some papers investigating the asymptotic error of score-based diffusion models, as in (Chen et al., 2023; Zhang et al., 2024).
@@
 These methods have relied heavily on various score-based samplers such as ULA or the reverse-time SDE of a diffusion process (Song et al., 2021).
-Notably, Sjöberg et al.
-(2023) describe a line integration method to compute the acceptance probabilities.
-Metropolis-Hastings Methods.
-Metropolis-Hastings (MH) algorithms are a class of MCMC algorithms that use an acceptance function to sample from target distributions (Metropolis et al., 1953; Hastings, 1970).
+Metropolis-Hastings.
+Metropolis-Hastings algorithms are a class of MCMC algorithms that use an acceptance function to sample from target distributions (Metropolis et al., 1953; Hastings, 1970).
 These algorithms generate a sequence of samples by proposing a candidate state from a proposal distribution and accepting or rejecting the candidate based on an acceptance criterion that ensures convergence to the target distribution.
 The efficiency of MH algorithms heavily depends on the choice of the proposal distribution, and significant research effort has focused on designing efficient proposal mechanisms (Rosenthal et al., 2011; Song et al., 2017; Titsias and Dellaportas, 2019; Davies et al., 2023; Lew et al., 2023).
-Recent developments in MH algorithms have explored the use of adaptive and state-dependent proposals to improve convergence rates and reduce autocorrelation in the generated samples (Andrieu and Thoms, 2008; Haario et al., 2001).
+Recent developments in MH algorithms have explored the use of adaptive proposals to improve convergence rates and reduce autocorrelation in the generated samples (Andrieu and Thoms, 2008; Haario et al., 2001).
 These adaptive MH methods adjust the proposal distribution during the sampling process to better capture the geometry of the target distribution, leading to faster and more accurate sampling (Roberts and Rosenthal, 2009; Brooks et al., 2011; Hirt et al., 2021; Biron-Lattes et al., 2024).
-(a) Original, weights $(0.8,0.2)$.
+(a) Original samples, weights $(0.8,0.2)$.
 (b) ULA samples, weights $(0.52,0.48)$.
@@
 3 Score-Based Metropolis-Hastings
+In this section, we first review score matching and MH algorithms in Sections 3.1 and 3.2 before we introduce Score Balance Matching in Section 3.3, a novel approach for learning the acceptance function from the score function.
 Let $X\sim p$ be a random variable, with support $\mathrm{supp}(X)=\mathcal{X}\subset\mathbb{R}^{d}$ and $p$ be its probability density function.
@@
 Denote these observed samples by $\{x^{\left(i\right)}\}_{i=1}^{N}$, where $N$ is the number of observed samples.
-The MH algorithm involves an acceptance function $a:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$, which defines the probability $a(x^{\prime},x)$ of transitioning from the current $x\in\mathcal{X}$ to a proposal $x^{\prime}\in\mathcal{X}$.
+3.1 Score Matching
+Score-based models are grounded in score matching (Hyvärinen, 2005), which minimizes the Fisher divergence between the estimated and true gradients of the log data distribution, defined as follows.
+Definition 1 (Fisher Divergence).
+The Fisher divergence $D_{\nabla}(p_{1}\|p_{2})$ between two probability distributions $p_{1}$ and $p_{2}$ is defined as:
+$D_{\nabla}(p_{1}\|p_{2})=\mathbb{E}_{x\sim p_{1}}\left[\|\nabla_{x}\log p_{1}(x)-\nabla_{x}\log p_{2}(x)\|^{2}\right].$
+Therefore, given a class of hypothesis function $\mathcal{S}\subset\{s:\mathcal{X}\rightarrow\mathbb{R}^{d}\}$, score matching aims to find $s^{*}\in\mathcal{S}$ that minimizes the Fisher divergence to the data distribution $p$:
+$s^{*}\in\operatorname*{arg\,min}_{s\in\mathcal{S}}\mathbb{E}_{x\sim p}\left[\|\nabla_{x}\log p(x)-s(x)\|^{2}\right]$ (1)
+Since the score function is not directly observed, Hyvärinen proved that minimizing in the loss funcition in Equation 1 is equivalent to minimizing the following loss function:
+$\mathbb{E}_{x\sim p}\left[\frac{1}{2}\|s(x)\|^{2}+\operatorname{tr}\left(\nabla_{x}s(x)\right)\right].$ (2)
+Several variants of this loss have been proposed to circumvent the cumbersome estimation of the Hessian term in Equation 2, including sliced score matching and denoising score matching, defined as follows:
+Definition 2 (Sliced Score Matching).
+Sliced score matching (SSM) projects the score function onto random directions $v\sim\mathcal{N}(0,I_{d})$:
+$\mathbb{E}_{x\sim p,v\sim\mathcal{N}(0,I_{d})}\left[\frac{1}{2}\|s(x)\|^{2}+v^{\top}\nabla_{x}s(x)v\right].$ (3)
+Definition 3 (Denoising Score Matching).
+Given a noise distribution $p_{\sigma}(\tilde{x}|x)$, the loss function is:
+$\mathbb{E}_{x\sim p,\tilde{x}\sim p_{\sigma}(\tilde{x}|x)}\left[\|\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x}|x)-s(\tilde{x})\|^{2}\right].$ (4)
+3.2 Metropolis-Hastings
+MH involves an acceptance function $a:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$, which defines the probability $a(x^{\prime},x)$ of transitioning from the current $x\in\mathcal{X}$ to a proposal $x^{\prime}\in\mathcal{X}$.
 The common choice of the acceptance function is given by:
@@
     • Draw $u\sim\text{Uniform}(0,1)$. If $u\leq a(x^{\prime},x_{t})$, accept $x_{t+1}=x^{\prime}$, otherwise set $x_{t+1}=x_{t}$.
-Note that for computing $a(x^{\prime},x_{t})$ in (1), a function proportional to $p(x)$ is needed.
-On the other hand, when sampling from $p(x)$, a common strategy is to estimate the score function $\nabla_{x}\log p(x)$ from data.
-In this work, we will assume that $\nabla_{x}\log p(x)$ is known through prior knowledge or estimated from the observed samples.
+Note that for computing $a(x^{\prime},x_{t})$ in Equation (5), a function proportional to $p(x)$ is needed.
 Detailed Balance.
-Although the acceptance function is often chosen according to (1), under mild conditions, the sufficient condition for convergence to the target distribution is that the acceptance ratio satisfies the detailed balance condition.
-The detailed balance condition, also known as the Markov reversibility condition, is defined as:
+Although the acceptance function is often defined by Equation (5), a sufficient condition for ensuring convergence to the target distribution is that the acceptance ratio satisfies the detailed balance condition:
 $\forall x,x^{\prime}\in\mathcal{X},\quad\frac{a(x^{\prime},x)}{a(x,x^{\prime})}=\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}.$ (6)
-Acceptance functions satisfying the detailed balance condition allow the Markov chain to have the target distribution as its unique stationary distribution under some mild assumptions.
-In particular, (1) is a particular solution to this condition.
-Our goal is then to use the knowledge of the stationary condition in (2) such that our sampling scheme is informed by this regime.
-Balance Matching.
-To do this, we develop a loss function to estimate a valid acceptance function from only the score function based on the detailed balance condition given in (2).
-In the following proposition, we present the main observation motivating our method.
-We denote by $\nabla$ the gradient with respect to $x,x^{\prime}\in\mathbb{R}^{d}$, i.e., $\nabla=\nabla_{x,x^{\prime}}$.
-Thus, for a function $f$ with values in $\mathbb{R}$, the gradient $\nabla f$ has values in $\mathbb{R}^{2d}$.
-Proposition 1.
-If the gradients of the log acceptance ratio satisfy the following condition: for every $x,x^{\prime}\in\mathcal{X}$,
-$\nabla\log\left(\frac{a(x^{\prime},x)}{a(x,x^{\prime})}\right)=\nabla\log\left(\frac{p(x^{\prime})}{p(x)}\right)+\nabla\log\left(\frac{q(x\mid x^{\prime})}{q(x^{\prime}\mid x)}\right)$ (3)
-then the acceptance function $a(x^{\prime},x)$ must satisfy the equality:
-$\frac{a(x,x^{\prime})}{a(x^{\prime},x)}=\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}.$
-Proof.
-Proof is provided in appendix C. ∎
-It follows from Proposition 1 that matching the gradients of the left-hand side and right-hand side in (2) is a sufficient condition for verifying the detailed balance condition. In other words, if:
-$\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})$
-$=\nabla\log p(x^{\prime})-\nabla\log p(x)+\nabla\log q(x\mid x^{\prime})$
-$\hskip 115.0pt-\nabla\log q(x^{\prime}\mid x),$
-then the detailed balance condition is satisfied.
-The acceptance function $a(x^{\prime},x)$ can be modeled using an expressive class of functions, such as neural networks, and this function can be optimized to satisfy the condition in (3).
-In this setup, the gradients $\nabla\log a(x^{\prime},x)$ and $\nabla\log a(x,x^{\prime})$ correspond to the model’s gradients, which can be efficiently computed via backpropagation when the model is differentiable.
-The terms $\nabla\log p(x)$ and $\nabla\log p(x^{\prime})$ represent the score functions at $x$ and $x^{\prime}$, which can be approximated with a score model.
-We can express these score functions as:
-$\nabla\log p(x)=\nabla_{x,x^{\prime}}\log p(x)=\left(\nabla_{x}\log p(x),\mathbf{0}\right)^{\top},$
-and, similarly,
-$\nabla\log p(x^{\prime})=\left(\mathbf{0},\nabla_{x^{\prime}}\log p(x^{\prime})\right)^{\top},$
-where $\mathbf{0}\in\mathbb{R}^{d}$ is the zero vector.
-Finally, we note that the gradients of the proposal distribution $q(x^{\prime}|x)$ are typically known by design when the proposal function is specified.
-MH with an Estimated Acceptance Functions.
-In cases where the only a score function is available, we propose to estimate the acceptance function with balance matching using the estimated score function so that MH algorithms can be applied.
-We term this approach Score-Based Metropolis-Hastings.
-For the scope of this paper, we will instantiate our method on three representative MH algorithms, specified by their proposal functions: the Random Walk Metropolis-Hastings (RW), the Metropolis-Adjusted Langevin Algorithm (MALA), and the Preconditioned Crank-Nicolson (pCN) algorithm.
-We next briefly review these methods.
-(a) Standard Random Walk Sampling.
-(b) Score-based RW with true scores.
-(c) Score-based RW with learned score.
-Figure 2: Comparison of sampling methods from a Gaussian distribution $\mathcal{N}(3,2)$ using three approaches: (a) Standard Random Walk Sampling with the acceptance function defined in (1), (b) Score-based RW with an acceptance network trained using true scores, and (c) Score-based RW with an acceptance network trained using a learned score.
-Random Walk Metropolis-Hastings: In the RW algorithm, the proposal function is typically a Gaussian centered at the current state, i.e., $q(x^{\prime}\mid x)=\mathcal{N}\left(x,\sigma^{2}I\right)$, where $\sigma^{2}$ controls the step size of the random walk.
-This proposal is simple and symmetric, making it easy to implement.
-Metropolis Adjusted Langevin Algorithm: MALA uses both the current state and the gradient of the log-probability.
-The proposal distribution is defined as $q(x^{\prime}\mid x)=\mathcal{N}\left(x+\frac{\varepsilon^{2}}{2}\nabla\log p(x),\varepsilon^{2}I\right)$, where $\varepsilon$ is a step size parameter.
-This approach leverages the local gradient information to propose moves that are more likely to be accepted.
-Preconditioned Crank-Nicolson: The pCN algorithm modifies the standard Crank-Nicolson approach by introducing a preconditioning step.
-The proposal function for pCN is given by $x^{\prime}=\sqrt{1-\beta^{2}}x+\beta\xi$, where $\xi\sim\mathcal{N}(0,I)$ is a standard Gaussian noise term and $\beta$ controls the step size.
-When the acceptance function is learned using balance matching, as defined in (3), we refer to these methods as Score RW, Score MALA, and Score pCN, respectively.
-While we illustrate our results on these three MH algorithms, we note that the proposed framework is general and can be applied to other proposal functions.
-Slow Mixing.
-There are two main motivations for our approach.
-The first is the challenge of applying traditional MH algorithms to generative modeling without requiring knowledge of the unnormalized density, while assuming that the score function can be accurately learned.
-The second motivation arises from limitations of the unadjusted Langevin algorithm (ULA), which, although commonly used for sampling from score-based models, does not guarantee convergence to the true distribution due to the impact of discretization.
-Consequently, by incorporating an MH adjustment step, we address a key limitation observed in the ULA when sampling from distributions with multiple modes, particularly when these modes are separated by regions of low density.
-Specifically, when the data distribution is a mixture $p(x)=\pi p_{1}(x)+(1-\pi)p_{2}(x)$, where $p_{1}(x)$ and $p_{2}(x)$ are distinct and largely disjoint distributions, the gradient of the log probability density, $\nabla_{x}\log p(x)$, becomes problematic.
-In regions where $p_{1}(x)$ dominates, the score is driven solely by $p_{1}(x)$, and similarly for $p_{2}(x)$.
-Consequently, ULA, which relies on these score gradients, fails to correctly sample from the mixture distribution as it does not properly account for the mixing proportions $\pi$ and $1-\pi$.
-This leads to an incorrect estimation of the relative densities between the modes.
-The faster mixing time of MALA compared to Langevin has been well-established in the literature as in (Dwivedi et al., 2019; Wu et al., 2022).
-To illustrate this, we conduct experiments to generate a mixture of Gaussian distributions with $p_{1}=\mathcal{N}\left((5,5)^{\top},I\right)$ and $p_{2}=\mathcal{N}\left((-5,-5)^{\top},I\right)$, with $\pi=0.8$.
-As can be observed in Figure 1, ULA produces samples that misrepresent the relative weights of the modes.
-On the other hand, we observe that the original MH algorithms significantly alleviate this problem, generating samples with mode weights close to the truth.
-This leads to our Score-Based MH framework, which facilitates the application of MH algorithms when only score functions are available.
+In particular, any acceptance function satisfying the detailed balance condition above ensures that the Markov chain has the target distribution $p(x)$ as its unique stationary distribution.
+Note that the standard acceptance function in Equation (5) is one specific solution that satisfies this condition.
 (a) $\lambda=0$
@@
 We plot the results for training an acceptance network with three different regularization values $\lambda\in\{0,0.1,1.0\}$
-4 Algorithm
-4.1 Acceptance Loss Function
-Ideal Acceptance Loss. We begin with the ideal scenario where the true score function $\nabla\log p$ is known. In this scenario, we propose the following loss function
-$\mathcal{L}_{i}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
+3.3 Score Balance Matching
+We now introduce our new objective function called Score Balance Matching (SBM).
+Let $q$ be a proposal distribution and let $\mathcal{A}$ be the set of valid acceptance functions for a given proposal $q$ and target $p$.
+[…]
+The SBM objective is defined as follows:
+$\mathcal{L}(a)$ $=\mathbb{E}_{x\sim p,x^{\prime}\sim q(\cdot|x)}\bigg{[}\left\|\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})\right.$
 $\quad-\left.\nabla\log p(x^{\prime})+\nabla\log p(x)\right.$
 $\quad\left.-\nabla\log q(x\mid x^{\prime})+\nabla\log q(x^{\prime}\mid x)\right\|^{2}\bigg{]},$ (7)
-with $X\sim p$ and $X^{\prime}\sim q$, and the optimization objective is to learn an acceptance function $a^{*}\in\mathcal{F}$, where $\mathcal{F}$ is a set of hypotheses, such that
-$a^{*}=\arg\min_{a\in\mathcal{F}}\mathcal{L}_{i}(a).$
+where $\nabla$ denotes the gradient with respect to $x,x^{\prime}\in\mathbb{R}^{d}$, i.e., $\nabla=\nabla_{x,x^{\prime}}$.
+The terms $\nabla\log p(x)$ and $\nabla\log p(x^{\prime})$ represent the score functions at $x$ and $x^{\prime}$, which can be approximated with a score model (e.g., obtained from score matching).
+We can express these score functions as:
+$\nabla\log p(x)=\nabla_{x,x^{\prime}}\log p(x)=\left(\nabla_{x}\log p(x),\mathbf{0}\right)^{\top},$
+and similarly,
+$\nabla\log p(x^{\prime})=\left(\mathbf{0},\nabla_{x^{\prime}}\log p(x^{\prime})\right)^{\top},$
+where $\mathbf{0}\in\mathbb{R}^{d}$ is the zero vector.
+Finally, we note that the gradients of the proposal distribution $q(x^{\prime}|x)$ are typically known by design when the proposal function is specified.
+Proposition 1.
+We have that, for every function $a:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$
+$\mathcal{L}(a)=0\iff a\in\mathcal{A}.$ (8)
+Therefore, if we minimize SBM to zero, we are guaranteed to have a valid acceptance function.
+In practice, the acceptance function $a$ can be parameterized using an expressive class of functions and optimized to minimize SBM.
+Since we estimate the acceptance function from finite samples, we establish a generalization bound to quantify the approximation error when learning $a(x^{\prime},x)$ using a function class $\mathcal{F}$ (e.g., neural networks).
+We start by making the following assumptions.
+Assumption 1.
+$\mathcal{X}$ is compact.
+Assumption 2.
+$\nabla\log p$ and $\nabla\log q$ are continuous on $\mathcal{X}\times\mathcal{X}$.
+Assumption 3.
+$\forall a\in\mathcal{F},\nabla\log a$ is continuous on $\mathcal{X}\times\mathcal{X}$.
+We note that Assumption 3 can be achieved by design. However, for many applications, Assumption 1 may not hold. To this end, we propose to employ the fact that the score function is scale invariant:
+$\nabla_{x}\log p(x)=\nabla_{x}\log p^{*}(x),\quad\forall x\in\mathcal{X}.$
+where, $\mathcal{X}\subset\mathbb{R}^{d}$ is chosen as a large compact subset of the support and $p^{*}$ is a truncated density:
+$p^{*}(x)=\frac{p(x)}{Z}\mathbbm{1}\left(x\in\mathcal{X}\right),$
+[…]
+Proposition 2 (Finite-Sample Generalization Bound).
+Let $\widehat{\mathcal{L}_{N}}(a)$ be the empirical SBM loss computed over $N$ i.i.d. training samples. Then, under Assumptions 12, and 3, we have that with probability at least $1-\delta$,
+$\sup_{a\in\mathcal{F}}\left|\mathcal{L}(a)-\widehat{\mathcal{L}_{N}}(a)\right|\leq 2\mathcal{R}_{N}(\mathcal{F})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{N}}\right),$ (9)
+where $\mathcal{R}_{N}(\mathcal{F})$ is the empirical Rademacher complexity of class $\mathcal{F}$.
+For the scope of this paper, we will instantiate our method on three representative MH algorithms, specified by their proposal functions: the Random Walk Metropolis-Hastings (RW), the Metropolis-Adjusted Langevin Algorithm (MALA), and the Preconditioned Crank-Nicolson (pCN) algorithm.
+We next briefly review these methods.
+RW: The proposal function for RW is a Gaussian centered at the current state, i.e., $q(x^{\prime}|x)=\mathcal{N}\left(x,\sigma^{2}I\right)$.
+MALA: The proposal distribution in MALA uses both the current state and the gradient of the log-probability, i.e., $q(x^{\prime}|x)=\mathcal{N}\left(x+\frac{\varepsilon^{2}}{2}\nabla\log p(x),\varepsilon^{2}I\right)$.
+pCN: The proposal function for pCN is given by $x^{\prime}=\sqrt{1-\beta^{2}}x+\beta\xi$, where $\xi\sim\mathcal{N}(0,I)$.
+When the acceptance function is learned using SBM, we refer to these methods as Score RW, Score MALA, and Score pCN, respectively.
+While we illustrate our results on these three MH algorithms, we note that the proposed framework is general and can be applied to other proposal functions.
+In Appendix A, we include another approach to approximate the acceptance function.
+(a) Original samples.
+(b) ULA.
+(c) Score RW.
+(d) Score MALA
+(e) Score pCN.
+Figure 3: Comparison of different methods on the Pinwheel dataset.
 (a) Original samples.
@@
 Figure 4: Comparison of different methods on the Moons dataset.
-(a) Original samples.
-(b) ULA.
-(c) Score RW.
-(d) Score MALA
-(e) Score pCN.
-Figure 5: Comparison of different methods on the Pinwheel dataset.
-Practical Acceptance Loss.
-Typically, only samples from the data-generating process are observed, and the true $\nabla\log p$ is unknown.
-However, score parameterization through a neural network has proven successful in practice.
-We assume that we have an approximation of the score function previously estimated through a model $\tilde{s}\left(x\right)\approx\nabla\log p\left(x\right)$.
-Note that this is often the case in practice where an estimated score function $\tilde{s}$ is used in conjunction with ULA for sampling.
-In this scenario, the proposed practical loss function is
-$\mathcal{L}_{p}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
-$\quad-\left.\tilde{s}(X^{\prime})+\tilde{s}(X)-\nabla\log q(X\mid X^{\prime})\right.$
-$\hskip 80.0pt\left.+\nabla\log q(X^{\prime}\mid X)\right\|^{2}\bigg{]},$ (5)
-with $X\sim p$ and $X^{\prime}\sim q$, and the optimization objective is to learn an acceptance function $\tilde{a}\in\mathcal{F}$ where
-$\tilde{a}=\arg\min_{a\in\mathcal{F}}\mathcal{L}_{p}(a).$
-Illustrative Example.
-We present the sampling results from a Gaussian distribution $\mathcal{N}(3,2)$ using the original Random Walk Metropolis-Hastings Algorithm, as shown in Figure 2(a).
-Next, we show the results of training with the ideal loss function for the Score-based RW in Figure 2(b), followed by the results using the practical loss function for Score-based RW in Figure 2(c).
-4.2 Entropy Regularization
-Moreover, our algorithm involves an entropy regularization term alongside the proposed loss function defined in Section 4.1.
-To illustrate its effectiveness, we present the following result as motivation.
-Proposition 2 (Estimated Acceptance Function).
-Let $x,x^{\prime}\in\mathcal{X}$, and let $M\geq 1$. Then, the acceptance function defined as:
-$a_{M}(x^{\prime},x)=\min\left\{\frac{1}{M},\frac{1}{M}\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}\right\}$
-satisfies the detailed balance condition in (2).
-Proof.
-Proof is provided in the appendix. ∎
-It follows from Proposition 2 that there are infinitely many solutions to the detailed balance condition, and the acceptance probabilities can become infinitesimally small.
+3.4 Motivation
+Our approach is motivated by three key challenges.
+First, traditional MH algorithms typically require the knowledge of the unnormalized density, while it is common in generative modeling to assume that only the score function can be accurately estimated.
+Second, ULA, widely used for sampling from score-based models, suffers from discretization errors that can lead to bias and slow-mixing.
+Consequently, by incorporating an MH adjustment step, we address a key limitation observed in ULA when sampling from distributions with multiple modes, particularly when these modes are separated by regions of low density.
+Specifically, when the data distribution is a mixture $p(x)=\pi p_{1}(x)+(1-\pi)p_{2}(x)$, where $p_{1}(x)$ and $p_{2}(x)$ are distinct and largely disjoint distributions, the gradient of the log probability density, $\nabla_{x}\log p(x)$, becomes misleading.
+In regions where $p_{1}(x)$ dominates, the score is driven solely by $p_{1}(x)$, and similarly for $p_{2}(x)$.
+Consequently, ULA, which relies on these score gradients, fails to correctly sample from the mixture distribution as it does not properly account for the mixing proportions $\pi$ and $1-\pi$.
+This leads to an incorrect estimation of the relative densities between the modes.
+The faster mixing time of MALA compared to Langevin has been well-established in the literature as in (Dwivedi et al., 2019; Wu et al., 2022).
+To illustrate this, we conduct experiments to generate a mixture of Gaussian distributions with $p_{1}=\mathcal{N}\left((5,5)^{\top},I\right)$ and $p_{2}=\mathcal{N}\left((-5,-5)^{\top},I\right)$, with $\pi=0.8$.
+Figure 1 shows that ULA produces samples that misrepresent the relative weights of the modes.
+In contrast, standard MH algorithms significantly alleviate this problem, generating samples with mode proportions that closely match the mixture weights.
+4 Algorithm
+In this section, we formulate the loss function and training algorithm for the acceptance function, which we model using neural networks.
+Entropy Regularization.
+In addition to the original objective function introduced in Equation (3.3), our algorithm involves an entropy regularization term.
+To illustrate its effectiveness, we present the following result as its motivation.
+Proposition 3.
+Let $a:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ and let $M\geq 1$ be a scaling factor. Then,
+$\mathcal{L}(a)=0\implies\mathcal{L}(\frac{a}{M})=0$
+It follows from Proposition 3 that there are infinitely many solutions to the detailed balance condition, including cases where the acceptance probabilities can become infinitesimally small.
 While each of these acceptance functions is a mathematically valid solution to the detailed balance condition, in practice, sampling with these low-acceptance probability functions can lead to always rejecting the proposal, resulting in a lack of convergence within a reasonable timeframe.
-Therefore, if we solely optimize the loss functions in (4.1) and (4.1), we risk converging to minima that are legitimate solutions but impractical due to their very low acceptance probabilities.
+Moreover, these infinitesimal values often cause training instabilities in the neural network that parameterizes the acceptance function.
+Therefore, if we solely optimize the loss function in (3.3), we risk converging to minima that are legitimate solutions but impractical due to their very low acceptance probabilities.
 To address this issue, we propose adding an entropy regularization term to our loss function.
 The modified loss function is given by:
-$\mathcal{L}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
+$\mathcal{L}_{r}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
 $-\tilde{s}(X^{\prime})+\tilde{s}(X)-\nabla\log q(X\mid X^{\prime})$
-$\;+\nabla\log q(X^{\prime}\mid X)\|^{2}\bigg{]}+\lambda\,\mathbb{E}\left[H(\log a(X^{\prime},X))\right],$ (6)
-where
+$\;+\nabla\log q(X^{\prime}\mid X)\|^{2}\bigg{]}+\lambda\,\mathbb{E}\left[H(a(X^{\prime},X))\right],$ (10)
+where:
 $H(a(X^{\prime},X))=$ $\log(a(X^{\prime},X))a(X^{\prime},X)$
 $+\log(1-a(X^{\prime},X))\left(1-a(X^{\prime},X)\right).$
-and $\lambda\geq 0$ is the weight of the regularization term.
-This regularization term encourages acceptance functions with higher entropy acceptance probabilities, thereby avoiding minima that result in very low acceptance probabilities.
+and $\lambda\geq 0$ is the weight of the regularization term. This regularization term encourages acceptance functions with higher entropy acceptance probabilities.
 (a) Original samples.
@@
 Figure 5: Comparison of different methods on the S-curve dataset.
-(a) Original samples.
-(b) ULA.
-(c) Score RW.
-(d) Score MALA.
-(e) Score pCN.
-Figure 7: Comparison of different methods on the Swiss Roll dataset.
 To illustrate the importance of the regularization term, we present the results using the Moons dataset from scikit-learn (Pedregosa et al., 2011).
@@
 Although this acceptance probability remains mathematically valid (i.e., it will still converge to the target distribution), the convergence time may become impractically long.
-4.3 Implementation
+Implementation.
+Typically, only samples from the target distribution $p$ are observed, and the true $\nabla\log p$ is unknown.
+However, score parameterization through a neural network has proven successful in practice.
+We assume that we have an approximation of the score function previously estimated through a model $\tilde{s}\left(x\right)\approx\nabla\log p\left(x\right)$.
+Note that this is often the case in practice where an estimated score function $\tilde{s}$ is used in conjunction with ULA for sampling.
 The final loss function used to train the acceptance network in our score MH framework consists of two components: the primary loss term, which enforces the reversibility condition, and an entropy regularization term.
@@
   • Using residual blocks akin to He et al. (2016).
-  • Employing smooth activation functions, such as GeLU or Softplus.
-  • Implementing annealed sampling from the proposal distribution, where instead of directly sampling from the target distribution, we start from a weighted combination of the observed data samples and the proposal samples.
+  • Employing smooth activation functions, e.g., GeLU.
+  • Clipping the gradients of the log acceptance function and the gradient of the log densities. This enables numerical stability and is also motivated by assumptions  2 and  3.
 Algorithm 1 Acceptance Loss Function
@@
 Output: Trained Acceptance Network $a$.
+(a) MALA, $\tau=0.5$
+(b) MALA, $\tau=0.7$
+(c) MALA, $\tau=0.9$
+(d) MALA, $\tau=1.1$
+(e) ULA, $\tau=0.5$
+(f) ULA, $\tau=0.7$
+(g) ULA, $\tau=0.9$
+(h) ULA, $\tau=1.1$
+Figure 6: MNIST: comparing annealed MALA (top) and annealed ULA (bottom) across $\tau$ values $0.5$, $0.7$, $0.9$, and $1.1$.
+Results show that MALA is more robust to step size.
+Both algorithms were run for $100$ denoising steps, with an additional $1000$ steps at the smallest noise level.
+$\tau$ is a fixed parameter that controls the adaptive step size.
 Table 1: Quantitative comparison of methods across the datasets using the following metrics: Wasserstein-1 (W1), Wasserstein-2 (W2), and Maximum Mean Discrepancy (MMD).
@@
 (a) W1 vs step size
-(b) W2 vs step size
-(c) MMD vs step size
 Figure 7: Influence of step size on the performance of Score-based MALA vs. ULA, evaluated on the Swiss Roll dataset.
 5 Empirical Results
-In this section, we present the empirical results on benchmark datasets including (Moons, Pinwheel, S-curve, and Swiss Roll), to demonstrate the validity of our approach.
-The results are presented in Figures 456, and 7.
+In this section, we present empirical results on various datasets (Moons, Pinwheel, S-curve, Swiss Roll, and MNIST), to demonstrate the validity of our approach.
+Further results on extreme value distributions are in Appendix B.
 Figure 8: Average acceptance ratio as a function of $\lambda$.
@@
 The results in Table 1 illustrate the performance of four sampling methods: ULA, Score-based RW, Score-based MALA, and Score-based pCN, across four datasets using three evaluation metrics: Wasserstein-1 (W1), Wasserstein-2 (W2), and Maximum Mean Discrepancy (MMD).
-Overall, the Score RW method consistently achieves lower W1 and W2 values compared to ULA across all datasets, highlighting its efficiency in sample-based inference.
+Figures 34, and  5 illustrate the results for low-dimensional data.
+Overall, score-based methods achieve lower W1 and W2 values compared to ULA across all datasets.
 This trend is especially noticeable on the Moons and Pinwheel datasets, where the W2 metric significantly outperforms that of the ULA method.
@@
 Stability with respect to Step Size: Score MALA vs. ULA.
-The results demonstrate that Score MALA is more robust to variations in step size compared to ULA.
-As seen in Figure 8, the W1, W2, and MMD metrics consistently show lower sensitivity to changes in step size for Score MALA, indicating its stability across different settings.
+The results demonstrate that Score MALA is more robust to variations in step size compared to ULA as seen in Figures 6,  7 and 9.
 Effect of Entropy Regularization.
@@
 This demonstrates that the performance remains robust for larger values of the entropy regularization term.
+Figure 9: MNIST data generation comparing annealed ULA and annealed score-based MALA. $\tau$ is a fixed parameter that controls the adaptive step size.
+Diffusion Models Application.
+We evaluate the performance of the the proposed method with denoising score matching applied to the MNIST dataset using annealed ULA and annealing with score-based MALA.
+Annealing  (Song and Ermon, 2019; Song et al., 2021) involves a sequence of noise scales, progressively refining the samples from high noise to low noise, thus improving the stability and efficiency of score-based sampling.
+Further details on the annealing process and its implementation are provided in the appendix.
+Our results presented in Figures 6 and  9 indicate that the proposed method exhibits greater stability with respect to step size variations compared to annealed ULA, as measured by the sensitivity parameter $\tau$.
+We present further details in Appendix E.
 6 Discussion
@@
 The main limitation of the method is the additional computational cost required to train the acceptance network.
-Furthermore, more advanced architectures need to be proposed to extend this framework to high-dimensional data.
+Furthermore, more advanced architectures need to be proposed to extend this framework to higher-resolution image data.
 Additionally, further theoretical analysis is needed to provide more appropriate priors on which acceptance function should be chosen.
 Acknowledgments
-Vahid Tarokh’s work was supported in part by the Air Force Office of Scientific Research under award number FA9550-20-1-0397.
+Ahmed Aloui and Vahid Tarokh’s work was supported in part by the Air Force Office of Scientific Research under award number FA9550-20-1-0397.
 Appendix
@@
 Third, we provide the proofs for the theoretical results.
-Finally, we present additional details of the experiments.
+Finally, we present additional details of the experiments and further empirical results.
 Appendix A Taylor Score-based Metropolis-Hastings
@@
 Proof of Proposition 1.
-Let $x,x^{\prime}\in\mathcal{X}$. Suppose that the gradients on both sides are equal, i.e.,
+We aim to prove that
+$\mathcal{L}(a)=0\iff a\in\mathcal{A}.$
+(Reverse Direction:) Suppose $a\in\mathcal{A}$, meaning that for every $x,x^{\prime}\in\mathcal{X}$, $a(x^{\prime},x)$ satisfies
+$\frac{a(x^{\prime},x)}{a(x,x^{\prime})}=\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}.$
+Taking the logarithm on both sides,
+$\log a(x^{\prime},x)-\log a(x,x^{\prime})=\log p(x^{\prime})-\log p(x)+\log q(x\mid x^{\prime})-\log q(x^{\prime}\mid x).$
+Differentiating both sides with respect to $x$ and $x^{\prime}$, we obtain
+$\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})=\nabla\log p(x^{\prime})-\nabla\log p(x)+\nabla\log q(x\mid x^{\prime})-\nabla\log q(x^{\prime}\mid x).$
+as we have that:
+$\mathcal{L}(a)=\mathbb{E}_{x\sim p,x^{\prime}\sim q(\cdot|x)}\bigg{[}\left\|\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})-\nabla\log p(x^{\prime})+\nabla\log p(x)-\nabla\log q(x|x^{\prime})+\nabla\log q(x^{\prime}|x)\right\|^{2}\bigg{]}$
+This implies $\mathcal{L}(a)=0$.
+(Forward Direction:)
+Suppose that $\mathcal{L}(a)=0$, i.e., as the expectation is taken over a positive variable we have that almost surely, for every $x,x^{\prime}\in\mathcal{X}$
 $\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})=\nabla\log p(x^{\prime})-\nabla\log p(x)+\nabla\log q(x\mid x^{\prime})-\nabla\log q(x^{\prime}\mid x),$
@@
 $1=e^{C}$
-Therefore we have that $C=0$, which concludes the proof. ∎
+Therefore we have that $C=0$. Hence, $a\in\mathcal{A}$, which concludes the proof. ∎
 Proof of Proposition 2.
+Let ^$\hat{\mathcal{L}}(a)$ be the empirical SBM loss computed over $N$ i.i.d. training samples:
+^$\hat{\mathcal{L}}(a)=\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla\log a(x_{i}^{\prime},x_{i})-\nabla\log a(x_{i},x_{i}^{\prime})-\nabla\log p(x_{i}^{\prime})+\nabla\log p(x_{i})-\nabla\log q(x_{i}\mid x_{i}^{\prime})+\nabla\log q(x_{i}^{\prime}\mid x_{i})\right\|^{2}.$
+The true expected loss is
+$\mathcal{L}(a)=\mathbb{E}_{x\sim p,x^{\prime}\sim q(\cdot|x)}\left[\left\|\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})-\nabla\log p(x^{\prime})+\nabla\log p(x)-\nabla\log q(x\mid x^{\prime})+\nabla\log q(x^{\prime}\mid x)\right\|^{2}\right].$
+All we need to prove is that the loss function is bounded in order to apply the standard uniform convergence bound from statistical learning theory (Mohri, 2018).
+Proving the boundedness of the loss function.
+By Assumption 1, $\mathcal{X}$ is compact, which implies that $\mathcal{X}\times\mathcal{X}$ is also compact.
+Furthermore, by Assumptions 2 and 3, the gradient terms $\nabla\log a$, $\nabla\log p$, and $\nabla\log q$ are continuous over a compact domain.
+By the Extreme Value Theorem, they attain their maximum absolute values, denoted by constants $C_{1},C_{2},$ and $C_{3}$, respectively:
+$\sup_{x,x^{\prime}\in\mathcal{X}}\|\nabla\log a(x^{\prime},x)\|\leq C_{1},\quad\sup_{x,x^{\prime}\in\mathcal{X}}\|\nabla\log p(x)\|\leq C_{2},\quad\sup_{x,x^{\prime}\in\mathcal{X}}\|\nabla\log q(x^{\prime}\mid x)\|\leq C_{3}.$
+Thus, the SBM loss is bounded as follows:
+$\mathcal{L}(a)\leq\mathbb{E}_{x\sim p,x^{\prime}\sim q(\cdot|x)}\left[(2C_{1}+2C_{2}+2C_{3})^{2}\right]=C_{4},$
+where $C_{4}=(2C_{1}+2C_{2}+2C_{3})^{2}$ is a finite constant.
+Applying standard uniform convergence bounds (Mohri, 2018), for all $a\in\mathcal{F}$, with probability at least $1-\delta$, we obtain
+$\sup_{a\in\mathcal{F}}\left|\mathcal{L}(a)-\hat{\mathcal{L}}(a)\right|\leq 2\mathcal{R}_{N}(\mathcal{F})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{N}}\right),$
+where $\mathcal{R}_{N}(\mathcal{F})$ is the Rademacher complexity of the function class $\mathcal{F}$. ∎
+Proof of Proposition 3.
+Suppose that $\mathcal{L}(a)=0$, Let $a_{M}=\frac{a}{M}$, we prove in Proposition 1 that $\mathcal{L}(a)=0$ is equivalent to $a\in\mathcal{A}$, hence we will prove that $a_{M}\in\mathcal{A}$.
 To prove that the acceptance function $a_{M}(x,x^{\prime})$ satisfies the detailed balance condition, we need to show that:
@@
 $\frac{1}{M}p(x^{\prime})q(x\mid x^{\prime})=\frac{1}{M}p(x^{\prime})q(x\mid x^{\prime}),$
-which confirms that the detailed balance condition holds in this case. ∎
+which confirms that the detailed balance condition holds in this case.
+Therefore, $a_{M}\in\mathcal{A}$. Hence, $\mathcal{L}(\frac{a}{M})=0$, which concludes the proof. ∎
 Appendix D Experiments Details
@@
 It is generated using sklearn.datasets.make_swiss_roll(n_samples=10000, noise=0.5).
+  • MNIST: MNIST consists of grayscale images of handwritten digits (0-9), each of size $28\times 28$.
+The dataset contains $60,000$ training samples and $10,000$ test samples.
+For our experiments, we normalize pixel values to $[-1,1]$ and convert the images into PyTorch tensors.
 All datasets are converted to PyTorch tensors for further processing. The experiments were performed using an NVIDIA RTX 3090 GPU.
@@
 Finally, the output is passed through a fully connected layer, followed by GELU and a Sigmoid activation to ensure the output is between 0 and 1, representing the acceptance probability.
+MNIST Architectures.
+For the MNIST Score Net, we employ a time-dependent UNet architecture with an encoder and decoder blocks and we use Softplus activations functions.
+For further details on the Score Network, please refer to the MNIST tutorial available at the following GitHub repository, as we use the same architecture https://github.com/yang-song/score_sde.
+For its acceptance network we use a Siamese like and that takes two inputs $(x,x^{\prime})$, embeds them via convolutional and residual layers, and processes them alongside learned time and step-size embeddings.
+We use GeLU activation functions.
+These branches do not share weights.
+The representation are then fed to a common branch.
+The acceptance network also encodes the time and step size parameters.
+Annealed ULA and MALA for MNIST.
+We evaluate annealed ULA and annealed MALA for sampling high-quality image reconstructions. We employ a denoisnig score matchign appraoch as presented in  (Song and Ermon, 2019; Song et al., 2021).
 D.3 Hyperparameters
@@
 Swiss Roll Adam $5\times 10^{-4}$ 512 4 200 1
+Appendix E Additional Empirical Results
+In this section, we include additional empirical results for three datasets.
+First, we perform a sanity check using a synthetic dataset, demonstrating that training the acceptance network with a learned score achieves similar performance to training with the true score, as shown in Figure 14.
+Furthermore, we present results for generating the Swiss Roll dataset in Figure 15.
+Finally, we provide additional results for the MNIST dataset, illustrating the robustness of Annealed MALA to the step size in Figure 16.
+(a) Standard Random Walk Sampling.
+(b) Score-based RW with true scores.
+(c) Score-based RW with learned score.
+Figure 14: Comparison of sampling methods from a Gaussian distribution $\mathcal{N}(3,2)$ using three approaches: (a) Standard Random Walk Sampling with the acceptance function defined in (5), (b) Score-based RW with an acceptance network trained using true scores, and (c) Score-based RW with an acceptance network trained using a learned score.
+(a) Original samples.
+(b) ULA.
+(c) Score RW.
+(d) Score MALA.
+(e) Score pCN.
+Figure 15: Comparison of different methods on the Swiss Roll dataset.
+Figure 16: Comparison of Wasserstein-1 (W1) values for Annealed MALA and Annealed ULA across different $\tau$ and step size settings.
+The plots classify regions based on whether the W1 values for both methods exceed or fall below a given threshold.
+Each subplot corresponds to a different threshold value ($30,32,34$).
+Colors indicate classification: (i) both methods above the threshold (blue), (ii) MALA below and ULA above (orange), (iii) ULA below and MALA above (green), and (iv) both methods below the threshold (gray).
+The x-axis represents the number of denoising steps, while the y-axis corresponds to the adaptive step size parameter $\tau$.
+We can see that annealed MALA exhibits a more robust behavior across different $\tau$ levels.